### v1

@@ 6 Numerical Experiments: Neural Network Approximations for High-Dimensional Problems
In this section, we provide a detailed analysis of the capability of neural networks to approximate solutions for high-dimensional problems with decaying sensitivity, as stated in Theorem 5.1. The first numerical experiment focuses on a LQR example, where the data of the optimal control problem exhibits a banded structure. In the second example, we consider a more general setting, in which the function is no longer quadratic but continues to display sensitivity decay. The numerical simulations reported in this paper are performed on an Nvidia GeForce RTX 3070. The codes are written in Python 3.9 with TensorFlow 2.10.
@@ 6.1 Linear quadratic optimal control problems
Consider a continuous-time linear quadratic optimal control problem as described in (13). The system matrix $A\in\mathbb{R}^{n\times n}$ is randomly generated with a bandwidth of 1, $i.e.$ the matrix is tridiagonal, and entries uniformly distributed in $(0,1)$, using the Python command numpy.random. Additionally, the matrices $B,Q$ and $R$ are all set to the identity matrix $I_{n}$. This induces a graph structure as in Figure 1 with $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\left\lvert{i-j}\right\rvert$. From Proposition 4.3 it follows that the solution $P$ of the correpsonding ARE (17) is an SED-matrix, which by Corollary 3.1 implies that $V(x)=x^{T}Px$ exhibits an exponentially decaying sensitivity. For dimension $n=200$, the decay of the first and hundredth row in $P$ are shown in the right column in Figure 4. To train the neural network to approximate the value function $V$ on the domain $\Omega\coloneqq[-1,1]^{200}$, we adopt a supervised learning approach with precomputed, randomly generated, and independently chosen training, validation and test data $\mathcal{D}_{\text{tr}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{te}}\subset\Omega$. These datasets are created by uniformly sampling data points from $\Omega$ and computing the corresponding values of $V$ using the solution $P$ of the ARE. The Mean Squared Error (MSE) on $\mathcal{D}\in\{\mathcal{D}_{\text{tr}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{te}}\}$ is given as
@@
Training is conducted on $\mathcal{D}_{\text{tr}}$ using a batch-size of $64$, the ADAM optimizer, and the sigmoid activation function $\sigma(x)=\frac{1}{1+e^{-x}}$. Every $10$ epochs, the MSE on the validation dataset $\mathcal{D}_{\text{val}}$ is evaluated to check whether it falls below a predefined tolerance $\varepsilon$. If this criterion is met or if the maximum number of epochs is reached, the training process is terminated. The final model is then verified by calculating the MSE on the test dataset.
@@
Table 1 compares training using the loss function $\nu_{g}=\nu_{z}=0.5$ with training based on the standard MSE loss, which corresponds to setting $\nu_{g}=\nu_{z}=0$ in $L$. The example is conducted using a fully connected neural network with a single hidden layer comprising 256 neurons across varying dimensions of the LQR problem. The network was trained with $|\mathcal{D}_{\text{tr}}|=|\mathcal{D}_{\text{val}}|=2^{17}$ data points until the MSE on the validation set reached a tolerance of $10^{-3}$. The results demonstrate that incorporating positive values for $\nu_{g}$ and $\nu_{z}$ significantly improves training efficiency, especially in high-dimensional settings. In particular, in the 50-dimensional case, setting $\nu_{g}=\nu_{z}=0.5$ allows the network to reach the tolerance after 40 epochs, whereas training with $\nu_{g}=\nu_{z}=0$ results in an MSE of $1.2\times 10^{-2}$ on the validation data after $1000$ epochs.
@@
10 Tolerance reached 10 Tolerance reached 60
20 Tolerance reached 20 Tolerance reached 200
50 Tolerance reached 40 Stopped after max. epochs 1000
@@ 20 Tolerance reached 20 Tolerance reached 200
Training with the loss function $L$ requires the computation of the gradient of the neural network with respect to the input, prior to evaluating the gradient of the loss with respect to the network parameters $\theta$. This is achieved using TensorFlow’s gradient tape, which enables automatic differentiation by recording operations on tensors during the forward pass and computing gradients during the backward pass. The benefits of such a modified loss function in the context of control systems have also been emphasized in the literature. For example, [2, 15, 1, 7] discuss the inclusion of the information on the derivative, while [14, 6] consider the incorporation of the loss at the origin.
@@
Figure 4 depicts the optimal value function $V$ and the trained neural network $W(x,\theta)$ projected to the $(x_{1},x_{2})-$ axis and the $(x_{100},x_{200})$ axis in the first two columns. The training was performed with $\nu_{g}=\nu_{z}=0.5$, $\left\lvert{\mathcal{D}_{\text{tr}}}\right\rvert=\left\lvert{\mathcal{D}_{\text{val}}}\right\rvert=131072$ points, and has reached a tolerance of $\varepsilon=10^{-3}$ after 110 epochs. The MSE, computed using $\left\lvert{\mathcal{D}_{\text{te}}}\right\rvert=2^{19}$ test points, is $9.1\cdot 10^{-4}$.
@@
From a theoretical perspective, any function that can be represented using the neural network architecture described in Figure 3 can also be represented by a standard feedforward neural network with one hidden layer. However, by explicitly encoding the separable structure within the network architecture, we observe a significant improvement in training performance compared to a fully connected feedforward network. Table 2 provides a brief comparison of a fully connected network and a S-NN for the $200$-dimensional LQR problem. When the matrix $A$ is fully connected, lacking any decaying property, the algorithm fails to converge for both network structures, with particularly poor performance observed for the S-NN. However, when $A$ has bandwidth $1$, introducing an exponential decaying sensitivity property, the fully connected network shows slightly improved performance but still fails to achieve the desired error tolerance within 1000 training epochs. In contrast, the S-NN meets the tolerance criterion $\varepsilon=10^{-3}$ after 110 epochs.
@@
200 2048 fully conn. - $1.7\times 10^{0}$ 1000
@@
1 2048 fully conn. - $4.7\times 10^{-1}$ 1000
1 16 S-NN 10 $9.1\times 10^{-4}$ 110
@@
Another criterion demonstrating the advantage of the S-NN architecture is the reduced number of training points and epochs required to achieve a test error below $\varepsilon=10^{-3}$. Consider a $50$ dimensional example where the matrix $A$ again has a bandwidth of $1$. We compare a fully connected single layer network with $13313$ parameters to a S-NN with $l=10$ and $9521$ parameters. Figure 6 presents the results obtained by varying the number of training points $|\mathcal{D}_{\text{tr}}|$. The maximum number of epochs was adjusted based on the size of the training dataset to ensure the algorithm had sufficient time to train: specifically, it was set to $1000$ for $2^{17}$ training points, $2000$ for $2^{16}$ points, and so on, doubling with each halving of the training set size. As before, the training loss parameters were set to $\nu_{g}=\nu_{z}=0.5$. We observe that, on one hand, the separable structured network requires significantly fewer training points to reach the desired tolerance in the test dataset, see Figure 6. On the other hand, for larger amounts of training data, where both architectures meet the tolerance, the S-NN requires considerably fewer epochs, see Figure 6.
@@
Furthermore, we compute approximations of the optimal value function $V$ for linear quadratic optimal control problems with increasing state dimensions, ranging from $20$ to $120$ in increments of $10$. Specifically, we considered two different cases: one where the matrix $A$ has bandwidth $1$, and another where $A$ has bandwidth $3$. We use a S-NN as shown in Figure 3 with $n=s$ subnetworks and $\nu_{g}=\nu_{z}=0.5$. The training dataset size was fixed at 4069 points, uniformly sampled in the unit ball, ensuring consistency with the framework established in Theorem 3.4. The graph-distance $l$ was fixed to $3$ in all cases. For each dimension, the size of the subsystems’ layers was minimized while maintaining a MSE in the test data below $10^{-4}$. It turned out that for both bandwidths and across all dimensions, only $M=2$ neurons per sublayer were required to satisfy this error threshold. Consequently, the total number of neurons and parameters increases linearly with the dimensionality of the problem, as illustrated in Figure 8. These findings suggest that the polynomial bound established in the theorems in Section 3 may not be sharp for LQR problems. It is important to note that, while the graph for the case with bandwidth equal to 1 is sequential, as shown in Figure 1, the increased number of connections in the case of bandwidth equal to 3 (depicted in Figure 8) leads to higher-dimensional graph neighborhoods. Specifically, for our choice of $l=3$ we have $d=4$ for the first case and $d=10$ for the second. This, in turn, results in a larger number of trainable parameters for the setting with bandwidth equal to $3$, as shown in Figure 8.
@@ 6.2 Approximation of a non-quadratic function
and consider a graph $\mathcal{G}$ with nodes $\{1,\dots,n\}$ and sequential connections leading to $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\left\lvert{i-j}\right\rvert$ for $i,j\in\{1,\dots,n\}$. Note that the function $V$ is not separable with respect to Definition 3.2 since each entry of its Hessian matrix is non-zero. However, for $\rho<1$ it exhibits an exponentially decaying sensitivity with basis $\rho$. We employ a S-NN with a distance for the graph-neighborhoods of $l=5$, a sublayer-size of $M=16$ in each subnetwork, $\nu_{g}=\nu_{z}=0.5$, and $8192$ training, validation, and test data. The resulting MSE is shown in Figure 10, where we observe that the function is not well approximated for $\rho=1$, corresponding to the case without decaying sensitivity. In contrast, for the case with exponentially decaying sensitivity, where $\rho=\frac{1}{8}$, the MSE falls below $10^{-4}$.

### v2

@@ 6 Numerical Experiments: Neural Network Approximations for High-Dimensional Problems
In this section, we provide a detailed analysis of the capability of neural networks to approximate solutions for high-dimensional problems with decaying sensitivity, as stated in Theorem 5.1. The first numerical experiment focuses on a LQR example, where the data of the optimal control problem exhibits a banded structure. In the second example, we consider a more general setting, in which the function is no longer quadratic but continues to display sensitivity decay. The numerical simulations reported in this paper are performed on an Nvidia GeForce RTX 3070. The codes are written in Python 3.9 with TensorFlow 2.10 and are available at https://github.com/MarioSperl/SNN-DecayingSensitivity.
@@ 6.1 Linear quadratic optimal control problems
Consider a continuous-time linear quadratic optimal control problem as described in (13). The system matrix $A\in\mathbb{R}^{n\times n}$ is randomly generated with a bandwidth of 1, $i.e.$ the matrix is tridiagonal, and entries uniformly distributed in $(0,1)$, using the Python command numpy.random. Additionally, the matrices $B,Q$ and $R$ are all set to the identity matrix $I_{n}$. This induces a graph structure as in Figure 2 with $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\left\lvert{i-j}\right\rvert$. From Proposition 4.3 it follows that the solution $P$ of the corresponding ARE (17) is an SED-matrix, which by Corollary 3.1 implies that $V(x)=x^{T}Px$ exhibits an exponentially decaying sensitivity. For dimension $n=200$, the decay of the first and hundredth row in $P$ are shown in the right column in Figure 5. To train the neural network to approximate the value function $V$ on the domain $\Omega=[-1,1]^{200}$, we adopt a supervised learning approach with precomputed, randomly generated, and independently chosen training, validation and test data $\mathcal{D}_{\text{tr}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{te}}\subset\Omega$. These datasets are created by uniformly sampling data points from $\Omega$ and computing the corresponding values of $V$ using the solution $P$ of the ARE. The Mean Squared Error (MSE) on $\mathcal{D}\in\{\mathcal{D}_{\text{tr}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{te}}\}$ is given as
@@
Training is conducted on $\mathcal{D}_{\text{tr}}$ using a batch-size of $64$, the ADAM optimizer with its default learning rate in TensorFlow, the sigmoid activation function $\sigma(x)=\frac{1}{1+e^{-x}}$, and TensorFlow’s default parameter initialization. These choices of hyperparameters for training have been applied consistently to all examples presented in this section. Every $10$ epochs, the MSE on the validation dataset $\mathcal{D}_{\text{val}}$ is evaluated to check whether it falls below a predefined tolerance $\varepsilon$. If this criterion is met or if the maximum number of epochs is reached, the training process is terminated. The final model is then verified by calculating the MSE on the test dataset.
@@
Table 1 compares training using the loss function $\nu_{g}=\nu_{z}=0.5$ with training based on the standard MSE loss, which corresponds to setting $\nu_{g}=\nu_{z}=0$ in $L$. The example is conducted using a fully connected neural network with a single hidden layer comprising 256 neurons across varying dimensions $n$ of the described LQR problem with $A$ having a bandwidth of $1$. The network was trained with $|\mathcal{D}_{\text{tr}}|=|\mathcal{D}_{\text{val}}|=2^{17}$ data points uniformly randomly sampled in $1,1]^{n}$ until the MSE on the validation set as defined in (23) reached a tolerance of $10^{-3}$. The results demonstrate that incorporating positive values for $\nu_{g}$ and $\nu_{z}$ significantly improves training efficiency, especially in high-dimensional settings. In particular, in the 50-dimensional case, setting $\nu_{g}=\nu_{z}=0.5$ allows the network to reach the tolerance after 240 epochs, whereas training with $\nu_{g}=\nu_{z}=0$ results in an MSE of $1.3\times 10^{-2}$ on the validation data after $10^{4}$ epochs.
@@
10 Tolerance reached 10 Tolerance reached 50
20 Tolerance reached 20 Tolerance reached 190
50 Tolerance reached 240 Stopped after max. epochs 1000
@@ 20 Tolerance reached 20 Tolerance reached 190
Training with the loss function $L$ requires the computation of the gradient of the neural network with respect to the input, prior to evaluating the gradient of the loss with respect to the network parameters $\theta$. This is achieved using TensorFlow’s gradient tape, which enables automatic differentiation by recording operations on tensors during the forward pass and computing gradients during the backward pass. The benefits of such a modified loss function in the context of control systems have also been emphasized in the literature. For example, [3, 19, 1, 10] discuss the inclusion of the information on the derivative, while [18, 7] consider the incorporation of the loss at the origin. Note that while the loss function $L$ is used for training, the mean squared error as defined in (23) is computed for the validation and test data.
@@
Figure 5 depicts the optimal value function $V$ and the trained neural network $W(x;\theta)$ projected to the $(x_{1},x_{2})$- axis and the $(x_{100},x_{200})$ axis in the first two columns. The training was performed with $\nu_{g}=\nu_{z}=0.5$, $\left\lvert{\mathcal{D}_{\text{tr}}}\right\rvert=\left\lvert{\mathcal{D}_{\text{val}}}\right\rvert=131072$ points, and has reached a tolerance of $\varepsilon=10^{-3}$ after 200 epochs. The MSE, computed using $\left\lvert{\mathcal{D}_{\text{te}}}\right\rvert=2^{19}$ test points, is ${5.4}\cdot 10^{-4}$.
@@
From a theoretical perspective, any function that can be represented using the neural network architecture described in Figure 4 can also be represented by a standard feedforward neural network with one hidden layer. However, by explicitly encoding the separable structure within the network architecture, we observe a significant improvement in training performance compared to a fully connected feedforward network. Table 2 provides a brief comparison of a fully connected network and a S-NN for the $200$-dimensional LQR problem using the same training parameters as in the experiment corresponding to Figure 5. When the matrix $A$ is fully connected, lacking any decaying property, the algorithm fails to converge for both network structures, with particularly poor performance observed for the S-NN. However, when $A$ has bandwidth $1$, introducing an exponential decaying sensitivity property, the fully connected network shows slightly improved performance but still fails to achieve the desired error tolerance within $10^{4}$ training epochs. In contrast, the S-NN meets the tolerance criterion $\varepsilon=10^{-3}$ after $200$ epochs.
@@
200 2048 fully conn. - ${1.9}\times 10^{0}$ 1000
@@
1 2048 fully conn. - ${3.5}\times 10^{-1}$ 1000
1 16 S-NN 10 ${5.4}\times 10^{-4}$ 200
@@
Another criterion demonstrating the advantage of the S-NN architecture is the reduced number of training points and epochs required to achieve a stopping tolerance of $\varepsilon=10^{-3}$. Consider a $50$ dimensional example where the matrix $A$ again has a bandwidth of $1$. We compare a fully connected single layer network with $13313$ parameters to a S-NN with $l=10$ and $9521$ parameters. Figure 7 presents the results obtained by varying the number of training points $|\mathcal{D}_{\text{tr}}|$ uniformly randomly chosen in $1,1]^{50}$. The maximum number of epochs was adjusted based on the size of the training dataset to ensure the algorithm had sufficient time to train: specifically, it was set to $1000$ for $2^{17}$ training points, $2000$ for $2^{16}$ points, and so on, doubling with each halving of the training set size. As before, the training loss parameters were set to $\nu_{g}=\nu_{z}=0.5$. We observe that, on one hand, the separable structured network requires significantly fewer training points to reach the desired tolerance in the test dataset, see Figure 7. On the other hand, for larger amounts of training data, where both architectures meet the tolerance, the S-NN requires considerably fewer epochs, see Figure 7.
@@
Furthermore, we compute approximations of the optimal value function $V$ for linear quadratic optimal control problems with increasing state dimensions, ranging from $20$ to $120$ in increments of $10$. Specifically, we considered two different cases: one where the matrix $A$ has bandwidth $1$, and another where $A$ has bandwidth $3$. We use a S-NN as shown in Figure 4 with $n=s$ subnetworks and $\nu_{g}=\nu_{z}=0.5$. The training dataset size was fixed at 4069 points, uniformly sampled in the unit ball, ensuring consistency with the framework established in Theorem 3.4. The graph-distance $l$ was fixed to $3$ in all cases. For each dimension, the size of the subsystems’ layers was minimized while maintaining an MSE in the test data below $10^{-4}$. It turned out that for both bandwidths and across all dimensions, only $M=2$ neurons per sublayer were required to satisfy this error threshold. Consequently, the total number of neurons and parameters increases linearly with the dimensionality of the problem, as illustrated in Figure 9. These findings suggest that the polynomial bound established in the theorems in Section 3 may not be sharp for LQR problems. It is important to note that, while the graph for the case with bandwidth equal to 1 is sequential, as shown in Figure 2, the increased number of connections in the case of bandwidth equal to 3 (depicted in Figure 9) leads to higher-dimensional graph neighborhoods. Specifically, for our choice of $l=3$ we have $d=4$ for the first case and $d=10$ for the second. This, in turn, results in a larger number of trainable parameters for the setting with bandwidth equal to $3$, as shown in Figure 9.
@@
Remark 6.1.
While the experiments in this section involving the S-NN network structure exploit the decaying sensitivity in the $P$-matrix to approximate $V(x)=x^{T}Px$, we would like to point out that this decay does not imply a decay in the singular values of $P$. This contrast is illustrated in Figure 10, which shows the decay of the first column of $P$ (left) and the singular values of $P$ (right), for a linear quadratic problem with tridiagonal $A\in\mathbb{R}^{n\times n}$ and $B=Q=R=I_{n}$, with $n=100$. We conclude that standard low-rank approximation techniques cannot exploit the decaying sensitivity property.
Figure 10: Comparison of the decaying sensitivity with the singular values of $P$.
@@ 6.2 Approximation of a non-quadratic function
and consider a graph $\mathcal{G}$ with nodes $\{1,\dots,n\}$ and sequential connections leading to $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\left\lvert{i-j}\right\rvert$ for $i,j\in\{1,\dots,n\}$. Note that the function $V$ is not separable with respect to Definition 3.2 since each entry of its Hessian matrix is non-zero. However, for $\rho<1$ it exhibits an exponentially decaying sensitivity with basis $\rho$. We employ a S-NN with a distance for the graph-neighborhoods of $l=5$, a sublayer-size of $M=16$ in each subnetwork, $\nu_{g}=\nu_{z}=0.5$, and $8192$ training, validation, and test data. The resulting MSE after training for $10^{4}$ epochs is shown in Figure 12, where we observe that the function is not well approximated for $\rho=1$, which corresponds to the case without decaying sensitivity. In contrast, for the case with exponentially decaying sensitivity, where $\rho=\frac{1}{8}$, the MSE falls below $10^{-4}$.
@@
6.3 Nonlinear optimal control problem
Finally, we turn our attention to a more challenging setting: the approximation of optimal value functions for systems governed by nonlinear dynamics. As a representative example, we consider the infinite-horizon optimal control problem associated with the Allen–Cahn equation, a nonlinear parabolic PDE of reaction-diffusion type, given by
$\left\{\begin{aligned} \partial_{t}y(t,x)&=\sigma\,\partial_{xx}y(t,x)+y(t,x)(1-y(t,x)^{2})+u(t,x),\\ y(0,x)&=y_{0}(x),\end{aligned}\right$.
on the spatial domain $x\in[0,1]$, equipped with homogeneous Neumann boundary conditions. The goal is to regulate the system toward the unstable steady state $\tilde{y}(x)\equiv 0$ by minimizing the quadratic cost
$\tilde{J}(y_{0},u)=\int_{0}^{\infty}\int_{0}^{1}\left(\delta_{1}|y(t,x)|^{2}+\delta_{2}|u(t,x)|^{2}\right)dx\,dt,\quad\delta_{1},\delta_{2}>0.$
The optimal value function for this nonlinear problem is not available in closed form. In order to be able to assess the performance of the separable neural network architecture for this example, we thus learn an approximation of the optimal value function to which we can then compare the function represented by the network. For obtaining this approximation, we discretize the state in space using finite differences with $n=50$ grid points, yielding a finite-dimensional nonlinear optimal control problem of the form
$\begin{split}\min_{u\in\mathcal{U}_{c}}J(u,y_{0})&=\int_{0}^{\infty}y(t)^{\top}Qy(t)+u(t)^{\top}Ru(t)\,dt,\\ \dot{y}(t)&=A(y(t))y(t)+Bu(t),\end{split}$ (24)
where $y(t)\in\mathbb{R}^{n}$, $A(y)\in\mathbb{R}^{n\times n}$ depends nonlinearly on the state, $B=I_{n}$, $Q=\frac{\delta_{1}}{n}I_{n}$, and $R=\frac{\delta_{2}}{n}I_{n}$. We approximate the optimal value function $V$ of (24) by a quadratic state-dependent ansatz: We set $V(y)=y^{\top}P(y)y$, where $P(y)$ is obtained as the solution to the State-Dependent Riccati Equation (SDRE):
$A^{\top}(y)P(y)+P(y)A(y)-P(y)BR^{-1}B^{\top}P(y)+Q=0.$ (25)
This construction does not solve the associated Hamilton–Jacobi–Bellman equation exactly, but provides a pointwise approximation of the value function by enforcing local optimality at each state $y$. The SDRE approach constitutes a generalization of the classical LQR scheme, designed to incorporate the state-dependent nonlinearities present in the system dynamics. or a detailed discussion of the theoretical underpinnings, we refer to [8] and the references therein, while applications in the context of optimal control for PDEs can be found in [1, 2]. Note that while $A(y)\in\mathbb{R}^{n\times n}$ depends on the state in this non-linear setting, its non-zero entries are independent of $y$. Thus, we can define the graph corresponding to $V$ with subsystems $z_{j}=y_{j}$, $1\leq j\leq n$, analogously to the linear quadratic setting described in Section 4 leading to a problem formulation fitting Setting 2.1 with $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=|i-j|$.
We now examine the emergence of a decaying sensitivity property in this nonlinear $50$-dimensional setting. By Definition 2.1 this requires to investigate the Lipschitz properties of the functions $g_{j}$ defined in (1). To this end, we fix the parameters $\sigma=10^{-2}$, $\delta_{1}=10$, $\delta_{2}=0.1$, and consider a fixed reference point $x\in\mathbb{R}^{n}$. For each index $i\in\{1,\ldots,n\}$, we define a perturbed vector $\tilde{x}_{i}\in\mathbb{R}^{n}$ by
$\tilde{x}_{i,k}=\begin{cases}0,&\text{if }k=i,\\ x_{k},&\text{otherwise}.\end{cases}$
A numerical estimate of the local Lipschitz constant of $g_{j}$ with respect to the $i$th coordinate is then given by
$\tilde{L}_{i,j}=\frac{\lvert g_{j}(x)-g_{j}(\tilde{x}_{i})\rvert}{\|x-\tilde{x}_{i}\|}.$
The right panels of Figure 13 report the empirical estimates $\tilde{L}_{i,j}$ for $j=1$ (top) and $j=25$ (bottom). Each estimate is computed at two reference points obtained by setting the $j$-th component to $\alpha$ and all other components to $0.1$, with $\alpha\in\{0.1,0.01\}$. Results are shown on a logarithmic scale. The observed exponential decay of $\tilde{L}_{i,j}$ with increasing |ij||i-j|, as well as its decrease for smaller $\|z_{j}\|$, provide strong numerical evidence that $V$ satisfies the decaying sensitivity property as defined in Definition 2.1.
This structural property justifies the use of the separable neural network architecture presented in Section 5 to leverage localized interactions to yield an accurate and efficient approximation of $V$. We employ a S-NN architecture configured with a graph‑neighborhood radius of $l=3$ and $M=32$ neurons in each subnetwork, resulting in a total of $9409$ trainable parameters. The dataset comprises $1024$ samples for training, validation, and testing uniformly randomly chosen in $\Omega=\{x\in\mathbb{R}^{50}\mid\|x\|_{2}\leq 1\}$, respectively. As regularization weights we chose $\nu_{g}=10^{4}$ and $\nu_{z}=10$. The large value of $\nu_{g}$ discourages flat regions in the surrogate, while the stringent tolerance guarantees a close fit to the reference solution; a systematic study of these effects is left for future work.
The left and middle panels in Figure 13 shows the results of a training process conducted with a stopping tolerance of $\epsilon=10^{-7}$. Note that while training was carried out on $\Omega$, i.e., the unit ball in $\mathbb{R}^{50}$, the plots are generated for visualization purposes on two-dimensional slices of the hypercube $1,1]^{50}$. We depict two representative coordinate planes, namely $(x_{1},x_{2})$ (top) and $(x_{25},x_{50})$ (bottom). The left panel depicts the true solution derived from the SDRE ansatz, while the middle panel shows the corresponding approximation generated by the S-NN. The visual agreement between the reference and surrogate value functions is strong, with the SNN model achieving a mean squared error in the test data of approximately $8.32\times 10^{-8}$ after 210 training epochs.
Figure 13: Plot of $V$ projected onto the $(x_{1},x_{2})$-axis (upper left) and onto the $(x_{25},x_{50})$-axis (lower left), the S-NN approximation onto the same hyperplanes (middle column), and the decay of $\tilde{L}_{i,j}$ for $j=1$ (upper right) and $j=25$ (lower right) varying the index ii.